# When Deepfakes Look Real: Detecting AI-Generated Faces with Unlabeled Data due to Annotation Challenges

### v2

@@ Introduction
$(\lambda_{\text{real}}^{i},\lambda_{\text{fake}}^{i})=h(z_{t}^{f}),\quad\lambda_{i}=\max(\lambda_{\text{real}}^{i},\lambda_{\text{fake}}^{i})$ (6)
where we retain samples of $\lambda_{i}\geq\lambda_{tf}$ to construct the feature library $\mathcal{B}$. $\mathcal{B}$ is split into real/fake sub-libraries $\mathcal{B}_{\text{real}}$ and $\mathcal{B}_{\text{fake}}$ according to the pseudo labels, where $\mathcal{B}_{\text{real}}$ and $\mathcal{B}_{\text{fake}}$ contain features labeled as real and fake according to the initial visual encoder predictions, respectively, which we use as a ‘simple’ reference case for curriculum learning.
Dynamic Threshold for Pseudo Label Generation. For each target sample $x_{t}^{j}$, we generate pseudo labels through a dual-verification process. First, consistent with the calculation in the feature library construction, we get the CLIP-based prediction $\hat{y}_{\text{clip}}^{j}=\lambda_{i}$. Next, we assess the feature library distance by calculating the minimum L2 distance between $z_{t}^{f}$ and the features in the real and fake sub-libraries:
$\quad d_{\text{fake}}^{j}=\min_{\mathbf{z}\in\mathcal{B}_{\text{fake}}}\|\mathbf{z}-z_{t}^{f}\|_{2}$ (7)
where we assign $\hat{y}_{\text{bank}}^{j}=0$ if $d_{\text{fake}}^{j}>0.5$, or $\hat{y}_{\text{bank}}^{j}=1$ otherwise. A pseudo label $\hat{y}_{unlab}^{j}$ is accepted if $\hat{y}_{\text{clip}}^{j}=\hat{y}_{\text{bank}}^{j}$ and $\hat{y}_{\text{clip}}^{j}\leq\lambda_{lt}^{(t)}$, where $\lambda_{lt}^{(t)}$ is a dynamic threshold. Based on the analysis of simple samples in the feature pool, we initialize $\lambda_{lt}^{(0)}=0.85$, which gradually decreases to 0.70 during training. This curriculum strategy initially prioritizes easier samples and progressively incorporates more challenging samples as $\lambda_{lt}^{(t)}$ decreases, ensuring robust learning from diverse target features $\mathcal{F}_{t}=\{z_{t}^{f}\mid x_{t}\in\mathcal{D}_{t}\}$ while minimizing bias toward less informative samples.
End-to-End Training Strategy
@@
End-to-End Loss Design. The overall training goal integrates the learning of the source domain and the target domain. The first stage is source domain training, and the second stage is joint training. The whole process is end-to-end, formulated as follows:
@@
Experiments
Settings
Datasets.
We used the following nine datasets: FaceForensics++ (FF++)(Rossler et al. 2019a), Deepfake Detection Challenge(DFDC)(Dolhansky et al. 2020a), preview version of DFDC(DFDCP)(Dolhansky et al. 2019), two versions of CelebDF (CDF-v1, CDF-v2)(Li et al. 2020c, b), DeepfakeDetection (DFD)(Deepfakedetection 2021), DF40(Yan et al. 2024c), UCDDP and UDF40, respectively; UCDDP is an unlabeled dataset obtained by sampling a small amount from the training sets of Deepfake Detection Challenge(DFDC), preview version of DFDC (DFDCP), two versions of CelebDF (CDF-v1, CDF-v2), DeepfakeDetection (DFD), and UDF40 is an unlabeled dataset obtained by sampling a small amount from the data subsets of different counterfeiting methods of DF40.
@@ Methods
FF++ (Rossler et al. 2019b) 1000 4000 5000 4 320k
CelebDF-v1 (Li et al. 2020b) $\dagger$ 408 795 1203 1 77k
CelebDF-v2 (Li et al. 2020b) $\dagger$ 590 5639 6229 1 399k
DFDCP (Dolhansky et al. 2019) $\dagger$ 1131 4119 5250 2 336k
DFDC (Dolhansky et al. 2020b) $\dagger$ 23654 104500 128154 8 8202k
DFD (Deepfakedetection 2021) $\dagger$ 363 3000 3363 5 215k
DF40 (Yan et al. 2024c) * $\sim$1500 0.1M+ 0.1M+ 40 1M+
UCDDP $\dagger$ 10+ 18k
UDF40 * 6 9k
@@
EfficientB4 (Tan and Le 2019) PMLR’19 EfficientNet 0.791 0.749 0.815 0.696 0.728 0.756
Xception(Chollet 2017) ICCV’19 Xception 0.779 0.737 0.816 0.708 0.737 0.755
Face X-ray (Li et al. 2020a) CVPR’20 HRNet 0.709 0.679 0.766 0.633 0.694 0.696
F3Net (Qian et al. 2020) AAAI’20 Xception 0.777 0.735 0.798 0.702 0.735 0.749
FFD (Dang et al. 2020) CVPR’20 Xception 0.784 0.744 0.802 0.703 0.743 0.755
SRM (Luo et al. 2021) CVPR’21 Xception 0.793 0.755 0.812 0.700 0.741 0.760
SPSL (Liu et al. 2021b) CVPR’21 Xception 0.815 0.765 0.812 0.704 0.741 0.767
Recce (Cao et al. 2022) CVPR’22 Designed 0.768 0.732 0.812 0.713 0.734 0.752
CORE (Ni et al. 2022) CVPR’22 Xception 0.780 0.743 0.802 0.705 0.734 0.753
UCF (Yan et al. 2023) ICCV’23 Xception 0.779 0.753 0.807 0.719 0.759 0.763
ED (Ba et al. 2024) AAAI’24 ResNet-34 0.818 0.864 - 0.721 0.851 -
LSDA (Yan et al. 2024a) CVPR’24 EfficientNet-B4 0.867 0.830 0.880 0.736 0.815 0.826
ProDet (Cheng et al. 2024) NeurIPS’24 EfficientNet-B4 0.909 0.844 - 0.811 0.724 0.822
UDD (Fu et al. 2025b) AAAI’25 ViT-B/16 - 0.869 0.910 0.758 0.856 -
Effort (Yan et al. 2024b) ICML’25 CLIP (ViT-L/14) 0.926 0.878 0.922 0.822 0.835 0.877
ForensicsAdapter (Cui et al. 2024) CVPR’25 CLIP (ViT-L/14) 0.914 0.900 0.933 0.843 0.890 0.896
@@
Table 2: Benchmark results for cross-dataset evaluation (Protocol‑1, frame-level AUC). All detectors are trained on FF++ c23(Rossler et al. 2019a) and evaluated on other deepfake datasets.
@@
F3Net (Qian et al. 2020) Xception 0.809 0.808 0.867 0.717 0.757 0.674 0.772
SPSL (Liu et al. 2021a) Xception 0.747 0.748 0.885 0.666 0.643 0.665 0.726
SRM (Luo et al. 2021) Xception 0.749 0.704 0.779 0.659 0.793 0.694 0.730
CORE (Ni et al. 2022) Xception 0.871 0.843 0.959 0.774 0.855 0.724 0.838
RECCE (Cao et al. 2022) Designed 0.898 0.832 0.925 0.848 0.848 0.768 0.853
SLADD (Chen et al. 2022) Xception 0.878 0.882 0.954 0.825 0.879 0.794 0.869
SBI (Shiohara and Yamasaki 2022) EfficientNet-B4 0.724 0.891 0.952 0.594 0.712 0.701 0.762
UCF (Yan et al. 2023) Xception 0.831 0.827 0.950 0.862 0.809 0.647 0.821
IID (Huang et al. 2023) Designed 0.839 0.789 0.888 0.844 0.789 0.644 0.799
LSDA (Yan et al. 2024a) EfficientNet-B4 0.872 0.875 0.930 0.721 0.855 0.793 0.841
ProDet (Cheng et al. 2024) EfficientNet-B4 0.908 0.929 0.975 0.747 0.837 0.844 0.873
CDFA (Lin et al. 2024b) SwinV2-B 0.762 0.756 0.823 0.803 0.772 0.757 0.779
ForensicsAdapter (Cui et al. 2024) CLIP (ViT-L/14) 0.969 0.886 0.963 0.943 0.937 0.917 0.936
Effort (Yan et al. 2024b) CLIP (ViT-L/14) 0.962 0.873 0.953 0.926 0.936 0.926 0.929
@@
Table 3: Benchmarking Results for Cross-Method Evaluation (Protocol-2, Video-Level AUC). All detectors are trained on FF++ c23 (Rossler et al. 2019b) and evaluated on other deepfake datasets.
Evaluation Protocol.
The results in Figure 2 show that simply using high-confidence pseudo labels to align with the domain does not lead to improved performance, especially for detectors with lower performance, and may even lead to a decrease in performance. Accordingly, we adopt two widely used standard protocols for evaluation: Protocol 1 is used for cross-dataset evaluation, and Protocol 2 is used for cross-operation evaluation within the FF++ domain. For Protocol 1, the model is trained using the labeled source domain dataset FF++ and the small unlabeled dataset UCDDP, and the performance is evaluated on the test set corresponding to the UCDDP sampling dataset (DFDC, DFDCP, CDF-v1, CDF-v2, DFD) to evaluate the generalization ability across datasets. For Protocol 2, the model is trained using FF++ and the small unlabeled dataset UDF40, and the performance is evaluated on the test subset corresponding to the UDF40 sampling dataset DF40 to evaluate the generalization ability across different forgery methods under a consistent data distribution. To further demonstrate that our performance improvement is not due to the use of unlabeled data, we let the baseline methods use the same unlabeled dataset (UCDDP or UDF40) for additional comparison by generating and utilizing pseudo labels.
Implementation Details.
We adopt CLIP ViT-L/14 (Radford et al. 2021) as the visual backbone, with input images resized to $\times224$ pixels. During training, we sample 16 frames per video, while 32 frames are used for testing. The model is optimized using the Adam optimizer (Kingma 2014) with a learning rate of 0.00008 and a weight decay of 0.0005. For training, the batch size is set to 32 for the source domain (FF++) and 10 for unlabeled data (UCDDP or UD40), with a test batch size of 32. Standard data augmentation techniques, including random cropping and flipping, are applied to enhance data diversity. For feature library construction, we set the initial confidence threshold to $\lambda_{tf}=0.9$. The dynamic pseudo-labeling threshold $\lambda_{\text{lt}}$ starts at 0.85 and gradually decreases to 0.70 during training. Loss hyperparameters $\lambda$, $\lambda_{1}$, $\lambda_{2}$, and $\beta$ are empirically set to 0.8, 0.4, 0.5, and 0.1, respectively. For evaluation, we report frame-level and video-level Area Under the Curve (AUC), a standard metric in deepfake detection, to compare our method with prior work. AUC provides a robust measure of classification performance across varying thresholds. All experiments are conducted on an NVIDIA RTX 4090 GPU.
@@
DANN 0.846 0.838 0.842
NAMC 0.878 0.877 0.878
SDAT 0.864 0.842 0.853
@@
Detection Performance
Table 2 presents the results of cross-dataset evaluation under Protocol-1. DPGNet achieves an average frame-level AUC of 0.938, surpassing the best baseline, ForensicsAdapter (0.896), by 4.2%. Notably, DPGNet excels on challenging datasets such as DFDC (AUC of 0.892, +4.9% over ForensicsAdapter) and CDF-v2 (AUC of 0.957, +5.7%). This superior performance is attributed to DPGNet’s text-guided alignment technique, which leverages text embeddings to capture domain-invariant forgery cues and effectively utilizes unlabeled data via a curriculum-driven pseudo labele generation technique. For Protocol-2, Table 3 reports cross-method evaluation results on DF40, where DPGNet achieves an average video-level AUC of 0.982, outperforming ForensicsAdapter (0.936) by 4.91%. DPGNet demonstrates significant improvements on advanced forgery methods, such as SimSwap (AUC of 0.984, +6.26%) and BleFace (AUC of 0.984, +5.92%), highlighting its robustness to diverse manipulation techniques within a consistent domain.
Figure 4: Cross-dataset (left) and cross-method evaluation.
Ablation Study
To dissect the contributions of the DPGNet design, we performed ablation studies on the unlabeled sample size, key components, loss functions, and unsupervised baselines, providing insights into its robust generalization capabilities across domains and forgery methods.
Core Components. Table 8 evaluates the core components of DPGNet: text-guided cross-domain alignment (TCA) and curriculum-driven pseudo label generation (CPG). Additionally, we provide a more in-depth evaluation of the cross-domain distillation (CD) strategy. The baseline without these components yields an AUC of 0.867. TGA alone boosts performance to 0.938 (+7.1%) by aligning visual-textual embeddings, mitigating domain gaps. Adding CPL increases the AUC to 0.946 (+7.9%) by progressively lowering the pseudo-labeling threshold to include challenging samples, while CD ensures robustness, maintaining the AUC at 0.953. This synergy drives its superior generalization.
Comparison with Unsupervised Methods. Table 10 compares DPGNet against unsupervised methods: DANN (Ganin and Lempitsky 2015), NAMC(Zhou et al. 2024), and source-free domain adaptation(SDAT). DPGNet outperforms these methods, utilizing text-guided alignment and curriculum learning to capture diverse forgery patterns while preserving source domain knowledge.
@@
Feature Distribution Visualization
To show the uniqueness of DPGNet from the feature distribution level, we use T-SNE (Van der Maaten and Hinton 2008) to visualize the feature distribution of the baseline and DPGNet. As shown in Figure 5, the baseline model exhibits significant overlap between real and fake features, indicating that it lacks semantic distinction in the target domain. In contrast, DPGNet learns more robust true/false semantics and significantly increases the separation between features of different categories. This larger separation surface DPGNet can more effectively bridge the gap between the source and target domains, thereby improving performance.
Conclusion
This work addresses the critical challenge of detecting deepfakes in realistic settings, where vast amounts of unlabeled data remain underutilized. We propose DPGNet, a novel framework that leverages text-guided alignment, curriculum-driven pseudo label generation, to fully exploit unlabeled deepfakes. By unifying visual and textual embeddings in a domain-invariant space, DPGNet captures generalizable features, dynamically selects informative samples to avoid overfitting, and preserves source-domain robustness via distillation. Extensive benchmarks (11 popular datasets) show that DPGNet consistently outperforms state-of-the-art methods (+6.3%). In the future, we plan to extend DPGNet with incremental learning and explore its applicability to related tasks such as anomaly detection and face anti-spoofing.

### v3

@@ 3.4 Curriculum-Driven Pseudo Label Generation
$(\lambda_{\text{real}}^{j},\lambda_{\text{fake}}^{j})=h(z_{t}^{f}),\quad\lambda_{j}=\max(\lambda_{\text{real}}^{j},\lambda_{\text{fake}}^{j})$ (6)
where we retain samples of $\lambda_{{j}}\geq\lambda_{tf}$ to construct the feature library $\mathcal{B}$. $\mathcal{B}$ is split into real/fake sub-libraries $\mathcal{B}_{\text{real}}$ and $\mathcal{B}_{\text{fake}}$ according to the pseudo labels, where $\mathcal{B}_{\text{real}}$ and $\mathcal{B}_{\text{fake}}$ contain features labeled as real and fake according to the initial visual encoder predictions, respectively, which we use as a ‘simple’ reference case for curriculum learning.
Dynamic Threshold for Pseudo Label Generation. For each target sample $x_{t}^{j}$, we generate pseudo labels through a dual-verification process. First, consistent with the calculation in the feature library construction, we get the CLIP-based prediction $\hat{y}_{{c}}^{j}=\arg\max(\lambda_{{real}}^{j},\lambda_{{fake}}^{j})$ and its confidence $\lambda_{j}=\max(\lambda_{{real}}^{j},\lambda_{{fake}}^{j})$. Next, we assess the feature library distance by calculating the minimum L2 distance between $z_{t}^{f}$ and the features in fake sub-libraries:
$d_{f}^{j}=\min_{\mathbf{z}\in\mathcal{B}_{{fake}}}\|\mathbf{z}-z_{t}^{f}\|_{2}$ (7)
where we assign $\hat{y}_{b}^{j}=0$ if $d_{f}^{j}>0.5$, or $\hat{y}_{b}^{j}=1$ otherwise. A pseudo label $\hat{y}_{unlab}^{j}$ is accepted if $\hat{y}_{c}^{j}=\hat{y}_{b}^{j}$ and the confidence $\lambda_{j}\geq\lambda_{lt}^{(t)}$, where $\lambda_{lt}^{(t)}$ is a dynamic threshold. Based on the analysis of simple samples in the feature pool, we initialize $\lambda_{lt}^{(0)}=0.85$, which gradually decreases to 0.70 during training. This curriculum strategy initially prioritizes easier samples and progressively incorporates more challenging samples as $\lambda_{lt}^{(t)}$ decreases, ensuring learning from diverse target features $\mathcal{F}_{t}=\{z_{t}^{f}\mid x_{t}\in\mathcal{D}_{t}\}$ while minimizing bias toward less informative samples.
3.5 End-to-End Training Strategy
@@
End-to-End Loss Design. The overall training goal integrates the learning of the source domain and the target domain. The first stage is source domain training, and the second stage is joint training. The whole process is end-to-end, and the specific training objectives are as follows:
@@
4 Experiments
4.1 Settings
Datasets. We used the following nine datasets: FaceForensics++(FF++) [ff++], Deepfake Detection Challenge(DFDC) [dfdc], preview version of DFDC(DFDCP) [dfdcp], two versions of CelebDF (CDF-v1, CDF-v2) [li2019celeb, li2020celeb], DeepfakeDetection (DFD) [dfd], DF40 [yan2024df40], UCDDP and UDF40, respectively; UCDDP is an unlabeled dataset obtained by sampling a small amount from the training sets of Deepfake Detection Challenge(DFDC), preview version of DFDC (DFDCP), two versions of CelebDF (CDF-v1, CDF-v2), DeepfakeDetection (DFD), and UDF40 is an unlabeled dataset obtained by sampling a small amount from the data subsets of different counterfeiting methods of DF40.
@@ Methods
FF++ [rossler2019faceforensics++] 1000 4000 5000 4 1:4 320k
CelebDF-v1 [li2020celeb] $\dagger$ 408 795 1203 1 $\sim$1:2 77k
CelebDF-v2 [li2020celeb] $\dagger$ 590 5639 6229 1 $\sim$1:10 399k
DFDCP [dfdcp] $\dagger$ 1131 4119 5250 2 $\sim$1:4 336k
DFDC [dolhansky2020deepfake] $\dagger$ 23654 104500 128154 8 $\sim$1:5 8202k
DFD [dfd] $\dagger$ 363 3000 3363 5 $\sim$1:10 215k
DF40 [yan2024df40] * $\sim$1500 0.1M+ 0.1M+ 40 $\sim$1:40 1M+
UCDDP $\dagger$ 10+ $\sim$1:5 18k
UDF40 * 6 $\sim$1:8 9k
@@
FFD [ffd] CVPR’20 Xception 0.784 0.744 0.802 0.703 0.743 0.755
SRM [luo2021generalizing] CVPR’21 Xception 0.793 0.755 0.812 0.700 0.741 0.760
SPSL [spsl] CVPR’21 Xception 0.815 0.765 0.812 0.704 0.741 0.767
RECCE [cao2022end] CVPR’22 Designed 0.768 0.732 0.812 0.713 0.734 0.752
CORE [ni2022core] CVPR’22 Xception 0.780 0.743 0.802 0.705 0.734 0.753
UCF [yan2023ucf] ICCV’23 Xception 0.779 0.753 0.807 0.719 0.759 0.763
[…]
LSDA [lsda] CVPR’24 EfficientNet-B4 0.867 0.830 0.880 0.736 0.815 0.826
[…]
Effort [effort] ICML’25 CLIP (ViT-L/14) 0.926 0.878 0.922 0.822 0.835 0.877
F-Adapter [ForensicsAdapter] CVPR’25 CLIP (ViT-L/14) 0.914 0.900 0.933 0.843 0.890 0.896
@@
Table 2: Benchmark results for cross-dataset evaluation (Protocol‑1, frame-level AUC). All detectors are trained on FF++ c23 [ff++] and evaluated on other deepfake datasets. $\dagger$ indicates that the result is taken from the original paper.
@@
SPSL [liu2021spatial] Xception 0.747 0.748 0.885 0.666 0.643 0.665 0.726
SRM [luo2021generalizing] Xception 0.749 0.704 0.779 0.659 0.793 0.694 0.730
CORE [ni2022core] Xception 0.871 0.843 0.959 0.774 0.855 0.724 0.838
RECCE [cao2022end] Designed 0.898 0.832 0.925 0.848 0.848 0.768 0.853
SLADD [chen2022self] Xception 0.878 0.882 0.954 0.825 0.879 0.794 0.869
SBI [shiohara2022detecting] EfficientNet-B4 0.724 0.891 0.952 0.594 0.712 0.701 0.762
UCF [yan2023ucf] Xception 0.831 0.827 0.950 0.862 0.809 0.647 0.821
IID [huang2023implicit] Designed 0.839 0.789 0.888 0.844 0.789 0.644 0.799
LSDA [lsda] EfficientNet-B4 0.872 0.875 0.930 0.721 0.855 0.793 0.841
ProDet [cheng2024can] EfficientNet-B4 0.908 0.929 0.975 0.747 0.837 0.844 0.873
CDFA [lin2024fake] SwinV2-B 0.762 0.756 0.823 0.803 0.772 0.757 0.779
F-Adapter [ForensicsAdapter] CLIP (ViT-L/14) 0.969 0.886 0.963 0.943 0.937 0.917 0.936
Effort [effort] CLIP (ViT-L/14) 0.962 0.873 0.953 0.926 0.936 0.926 0.929
@@
Table 3: Benchmarking Results for Cross-Method Evaluation (Protocol-2, Video-Level AUC). All detectors are trained on FF++ c23 [rossler2019faceforensics++] and evaluated on other deepfake datasets.
Evaluation Protocol.The results in Figure 2 show that simply using high-confidence pseudo labels to align with the domain does not lead to improved performance, especially for detectors with lower performance, a large number of pseudo-label generation errors can even lead to performance degradation. Accordingly, we adopt two widely used standard protocols for evaluation: Protocol 1 is used for cross-dataset evaluation, and Protocol 2 is used for cross-operation evaluation within the FF++ domain. For Protocol 1, the model is trained using the labeled source domain dataset FF++ and the small unlabeled dataset UCDDP, and the performance is evaluated on the test set corresponding to the UCDDP sampling dataset (DFDC, DFDCP, CDF-v1, CDF-v2, DFD) to evaluate the generalization ability across datasets. For Protocol 2, the model is trained using FF++ and the small unlabeled dataset UDF40, and the performance is evaluated on the test subset corresponding to the UDF40 sampling dataset DF40 to evaluate the generalization ability across different forgery methods under a consistent data distribution. To further demonstrate that our performance improvement is not due to the use of unlabeled data, we let the baseline methods use the same unlabeled dataset (UCDDP or UDF40) for additional comparison by generating and utilizing pseudo labels.
Implementation Details. We adopt CLIP ViT-L/14 [clip_paper] as the visual backbone, with input images resized to $\times224$ pixels. During training, we sample 16 frames per video, while 32 frames are used for testing. The model is optimized using the Adam optimizer [adam] with a learning rate of 0.00008 and a weight decay of 0.0005. For training, the batch size is set to 32 for the source domain (FF++) and 10 for unlabeled data (UCDDP or UD40), with a test batch size of 32. Standard data augmentation techniques, including random cropping and flipping, are applied to enhance data diversity. For feature library construction, we set the initial confidence threshold to $\lambda_{tf}=0.9$. The dynamic pseudo-labeling threshold $\lambda_{\text{lt}}$ starts at 0.85 and gradually decreases to 0.70 during training. Loss hyperparameters $\lambda$, $\lambda_{1}$, $\lambda_{2}$, and $\beta$ are empirically set to 0.8, 0.4, 0.5, and 0.1, respectively. For evaluation, we report frame-level and video-level Area Under the Curve (AUC), a standard metric in deepfake detection, to compare our method with prior work. AUC provides a robust measure of classification performance across varying thresholds. All experiments are conducted on an NVIDIA RTX 4090 GPU.
@@
4.2 Detection Performance
Table 2 presents the results of cross-dataset evaluation under Protocol-1. DPGNet achieves an average frame-level AUC of 0.938, surpassing the best baseline, ForensicsAdapter (0.896), by 4.2%. Notably, DPGNet excels on challenging datasets such as DFDC (AUC of 0.892, +4.9% over ForensicsAdapter) and CDF-v2 (AUC of 0.957, +5.7%). This superior performance is attributed to DPGNet’s text-guided alignment technique. For Protocol-2, Table 3 reports cross-method evaluation results on DF40, where DPGNet achieves an average video-level AUC of 0.982, outperforming ForensicsAdapter (0.936) by 4.91%. DPGNet demonstrates significant improvements on advanced forgery methods, such as SimSwap (+6.26%) and BleFace (+5.92%), highlighting its generalization to diverse manipulation techniques within a consistent domain.
4.3 Ablation Study
To dissect the contributions of the DPGNet design, we performed ablation studies on the unlabeled sample size, key components, loss functions, and unsupervised baselines.
Core Components. Table 8 evaluates the core components of DPGNet: text-guided cross-domain alignment (TCA) and curriculum-driven pseudo label generation (CPG). Additionally, we provide a more in-depth evaluation of the cross-domain distillation (CD) strategy. The baseline without these components yields an AUC of 0.867. TGA alone boosts performance to 0.938 by aligning visual-textual embeddings, mitigating domain gaps. Adding CPL increases the AUC to 0.946 by progressively lowering the pseudo-labeling threshold to include challenging samples, while CD ensures generalization, maintaining the AUC at 0.953. This synergy drives its superior generalization.
Comparison with Unsupervised Methods. Table 10 compares DPGNet against unsupervised methods: DANN [ganin2015unsupervised], NAMC [unsurpervised], and Source-free Domain Adaptation [li2024comprehensive]. DPGNet outperforms these methods, utilizing text-guided alignment and curriculum learning to capture diverse forgery patterns while preserving source knowledge.
@@ 4.3 Ablation Study
DANN[ganin2015unsupervised] 0.846 0.838 0.842
NAMC [unsurpervised] 0.878 0.877 0.878
SDAT [li2024comprehensive] 0.864 0.842 0.853
@@
4.4 Feature Distribution Visualization
To show the uniqueness of DPGNet from the feature distribution level, we use T-SNE [van2008visualizing] to visualize the feature distribution of the baseline and DPGNet. As shown in Figure 4, the baseline model exhibits significant overlap between real and fake features, indicating that it lacks semantic distinction in the target domain. In contrast, DPGNet learns more general true/false semantics and significantly increases the separation between features of different categories. This larger separation surface DPGNet can more effectively bridge the gap between the source and target domains, thereby improving performance.
5 Conclusion
This work addresses the critical challenge of detecting deepfakes in realistic settings, where vast amounts of unlabeled data remain underutilized. We propose DPGNet, a novel framework that leverages text-guided alignment, curriculum-driven pseudo label generation, to exploit unlabeled deepfakes. By unifying visual and textual embeddings in a domain-invariant space, DPGNet captures generalizable features, selects informative samples to avoid overfitting. Extensive benchmarks show that DPGNet outperforms SOTA methods by a margin.